OpenStudy (anonymous):

Stan’s savings account has a balance of $1986. After 23 years, what will the amount of interest be at 4% compounded annually? a. $2908.93 c. $2913.93 b. $2899.93 d. $794.40

OpenStudy (rane):

do u know the formula for finding compound interest ?

OpenStudy (anonymous):

nope

OpenStudy (johnweldon1993):

Do you got this @RANE ?

OpenStudy (anonymous):

I need to submit this in 10min so any help would be greatly appreciated!

OpenStudy (johnweldon1993):

Well @GMM You would use the equation \[P = C(1 + r)^t\] Where r = your rate *in a decimal form* C = your original deposit t = time *in years* So plugging in your numbers you get... \[P = 1986(1 + .04)^{23}\] so \[P = 1986(1.04)^{23}\] What do you get when you solve that?

OpenStudy (anonymous):

wow your fast johnweldon :)

OpenStudy (johnweldon1993):

Thank you @kelliegirl33 lol :)

OpenStudy (johnweldon1993):

@GMM *Hint ...remember to raise 1.04 to the 23rd power first! Then multiply that answer by 1986

OpenStudy (anonymous):

4706.65871993 RadDeg x! Inv sin ln π cos log e tan √ Ans EXP xy ( ) % AC 7 8 9 ÷ 4 5 6 × 1 2 3 - 0 . = +

OpenStudy (anonymous):

lol yeah Im thinking Im wrong

OpenStudy (rane):

@johnweldon1993 yes i did get it bt i wasnt here

OpenStudy (johnweldon1993):

Ahh okay...sorry for taking over @RANE :) @GMM \[P = 1986(1.04)^{23}\] I get \[P = 4894.92506871\] Now this is what the total is...you dont want that...you want JUST the interest earned....so take that number....and subtract 1986 from it to find your answer

OpenStudy (anonymous):

LOVE YOU thanks!!!!!!!!!!!

OpenStudy (rane):

its not a problem @johnweldon1993

OpenStudy (rane):

lol @GMM

OpenStudy (johnweldon1993):

lol No problem @GMM